# Spectrum and Resonances of Stark Hamiltonians with $\delta$–Shell Potentials

### v1

@@ Abstract
We study the Stark Hamiltonian with a $\delta$–interaction supported on a hypersphere, with boundary strength $\alpha\in L^{\infty}(S_{a})$. Using the framework of quasi boundary triples, we construct self–adjoint realizations and derive a Krein–type resolvent formula. For $F\neq 0$ we prove $\sigma(H_{F,\alpha})=\sigma_{\mathrm{ac}}(H_{F,\alpha})=\mathbb{R}$ by transferring a strict Mourre estimate from the free Stark operator and by the compactness of the boundary coupling. Resonances are defined by complex distortion in the field direction and characterized as zeros of a boundary regularized Fredholm determinant. In $d=3$ and small $F$, bound states at $F=0$ continue to resonances with widths of order $\exp(-c/F)$, governed by an Agmon–type distance. A brief comparison with the free Stark case is also given.
@@ 1 Introduction
    Small–field asymptotics in $d=3$. For small $F>0$, bound states at $F=0$ continue to resonances whose widths are exponentially small in $1/F$, governed by an Agmon–type distance.
@@ 2 Boundary triples and self–adjoint extensions
Here $\Gamma_{0}$ and $\Gamma_{1}$ denote the boundary maps defined above, that is, $\Gamma_{0}$ is the jump of the normal derivative across $S_{a}$ and $\Gamma_{1}$ is the average trace on $S_{a}$. For $z\in\mathbb{C}\setminus\mathbb{R}$ let $N_{z}=\ker(T^{*}-z)$. Then the map
$\Gamma_{0}:N_{z}\longrightarrow H^{-1/2}(S_{a})$
@@
Surjectivity follows from Lemma 3, by combining the interior and exterior solutions with a common trace $\phi$. If $u\in N_{z}$ satisfies $\Gamma_{0}u=0$, then $u^{\pm}$ solve homogeneous Dirichlet problems with nonreal energy, which implies $u^{\pm}\equiv 0$. Hence $\Gamma_{0}$ is also injective. The boundedness of the inverse map follows from the closed graph theorem. $\Box$
Lemma 4 shows that $\Gamma_{0}:N_{z}\to H^{-1/2}(S_{a})$ is bijective for every nonreal $z$. Since $H^{1/2}(S_{a})$ is infinite dimensional, it follows that the deficiency indices of $T$ are infinite, so $T$ is not essentially self–adjoint. Thus self–adjoint realizations arise precisely from boundary conditions relating $\Gamma_{0}$ and $\Gamma_{1}$. In particular, when the wall strength varies along the surface, $\alpha(\omega)\in L^{\infty}(S_{a};\mathbb{R})$, we introduce the boundary multiplication operator
@@ 4 Spectrum of the penetrable–wall Stark Hamiltonian
Step 3: Mourre estimate and absence of eigenvalues. Take $A_{F}:=\mathrm{sgn}(F)\,p_{1}$ with $p_{1}=-i\partial_{x_{1}}$. Then $\Delta,\,iA_{F}]=0$ and $Fx_{1},\,iA_{F}]=|F|$, hence $H_{F,0},\,iA_{F}]=|F|I$. By [15, Def. 2.1] the pair $(H_{F,\alpha},A_{F})$ is of class $C^{1}$ on any bounded interval: indeed $H_{F,0},iA_{F}]=|F|$ and the boundary perturbation $A_{\alpha}\delta(|x|-a)$ is $H_{F,0}$–compact by Lemma 8, so the commutator with $A_{F}$ is well defined in the form sense after spectral localization. Together with the compactness of $E_{J}(H_{F,\alpha})-E_{J}(H_{F,0})$ (Step 2), the abstract Mourre theory [15, Thm. 2.2 & Thm. 4.3] (see also [1, Chap. 7]) yields the transfer of the strict Mourre estimate from $H_{F,0}$ to $H_{F,\alpha}$ up to a compact remainder. In particular,
$E_{J}(H_{F,\alpha})[H_{F,\alpha},iA_{F}]E_{J}(H_{F,\alpha})=|F|\,E_{J}(H_{F,\alpha})+R_{J},\qquad\|R_{J}\|<|F|/2,$
where $R_{J}$ is compact (the bound holds after choosing $J$ sufficiently small). Here the contribution of the boundary interaction $A_{\alpha}\delta(|x|-a)$ yields only compact terms after localization. $\Box$
@@ 5.1 Complex distortion and analytic continuation
This defines a type–A analytic family of operators in the sense of Kato. The distorted resolvent
@@ 5.2 Distorted boundary operators and Krein formula
and is denoted by $H_{F,\alpha}(\theta)$. For $0<\text{{\rm Im}}\theta<\theta_{0}$, $U_{\theta}$ is boundedly invertible on $L^{2}$ and the distorted free operator $H_{F,0}(\theta)=U_{\theta}H_{F,0}U_{\theta}^{-1}$ is mm–sectorial [12, 13]. We keep the real boundary maps $(\Gamma_{0},\Gamma_{1})$ from Section 2, so the $\delta$–shell boundary condition $\Gamma_{0}=A_{\alpha}^{b}\Gamma_{1}$ remains unchanged under the distortion. What changes is only the Poisson and Weyl operators $\gamma_{F}(\theta;z)$ and $M_{F}(\theta;z)$, which are now constructed using the distorted free resolvent $R_{F,0}(\theta;z)$. Within the quasi boundary triple framework this leads to a closed mm–sectorial extension, whose resolvent is given by the distorted Kreĭn formula (see (5.2)); cf. [3, Sec. 6, Thm. 6.16] and [2, Thm. 2.6]. For $\theta\in\mathbb{R}$ we recover the undistorted self–adjoint extension $H_{F,\alpha}$.
@@
Step 3 (Ansatz and boundary equation). Let $f\in L^{2}(\mathbb{R}^{d})$ and $z\in\rho(A_{0}(\theta))\cap\rho(A_{\alpha}(\theta))$. Set $u_{0}=(A_{0}(\theta)-z)^{-1}f$, so $(T(\theta)^{*}-z)u_{0}=f$ and $\Gamma_{0}u_{0}=0$. Seek $u=u_{0}-\gamma_{F}(\theta;z)\varphi$ with unknown $\varphi\in H^{-1/2}(S_{a})$. Then $\Gamma_{0}u=-\varphi$, $\Gamma_{1}u=\Gamma_{1}u_{0}-M_{F}(\theta;z)\varphi$. Imposing the boundary condition $\Gamma_{0}=A_{\alpha}^{b}\Gamma_{1}$ yields
$\Gamma_{1}u_{0}=(M_{F}(\theta;z)-A_{\alpha}^{b})\varphi,$
hence (since $z\in\rho(A_{\alpha}(\theta))$)
$\varphi=-(I-A_{\alpha}^{b}M_{F}(\theta;z))^{-1}\Gamma_{1}u_{0}.$
Step 4 (Passage to the $L^{2}$ pivot). Note the similarity identity
$\Lambda^{-1}(I-A_{\alpha}^{b}M_{F}(\theta;z))\Lambda=I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z),$
whence
$(I-A_{\alpha}^{b}M_{F}(\theta;z))^{-1}=\Lambda(I-A_{\alpha}^{\flat}M_{F}^{\flat}(\theta;z))^{-1}\Lambda^{-1}.$
Inserting this into the ansatz for $u$ and recalling that $u=(H_{F,\alpha}(\theta)-z)^{-1}f$ yields (5.2), which completes the proof. $\Box$
@@ 5.3 Boundary characterization of resonances
By the Herbst distortion theory (Theorem in the previous subsection), if $0<\text{{\rm Im}}~\theta<\theta_{0}$, the discrete eigenvalues of $H_{F,\alpha}(\theta)$ in the lower half-plane are independent of $\theta$. Therefore, the set of resonances of $H_{F,\alpha}$ is well defined and coincides with the zero set of $D_{F}^{(p)}(\theta;z)$ for any admissible $\theta$.
@@ 5.4 Small–field asymptotics in three dimensions
We finally examine the fate of bound states under the application of a small Stark field. Throughout we assume $d=3$.
While the general boundary operator framework applies to $\delta$–interactions on smooth compact hypersurfaces, in this section we specialize to the case of a spherical shell. This restriction allows us to exploit spherical symmetry and obtain explicit formulas for bound states and their continuation to resonances.

### v2

@@ Abstract
We study the Stark Hamiltonian with a $\delta$–interaction supported on a hypersphere. Using the framework of quasi boundary triples, we construct self–adjoint realizations and derive a Krein–type resolvent formula. For $F\neq 0$ we show that the spectrum is purely absolutely continuous and equals the whole real line, by transferring a Mourre estimate from the free Stark operator. Resonances are defined by complex distortion and characterized as zeros of a boundary regularized Fredholm determinant. In three dimensions and for small $F$, bound states at $F=0$ continue to resonances: their energies admit a quadratic shift, and their widths are exponentially small in $1/F$, with rate determined by an Agmon action. A brief comparison with the free Stark case is also given.
@@ 1 Introduction
    Small–field asymptotics in $d=3$. For small $F>0$, bound states at $F=0$ continue to resonances. Their energies admit a perturbative expansion with a quadratic shift, and their widths are exponentially small in $1/F$, with the optimal decay rate determined by an Agmon action.
@@ 2 Boundary triples and self–adjoint extensions
Here $\Gamma_{0}$ and $\Gamma_{1}$ denote the boundary maps defined above, that is, $\Gamma_{0}$ is the jump of the normal derivative across $S_{a}$ and $\Gamma_{1}$ is the average trace on $S_{a}$. For $z\in\mathbb{C}\setminus\mathbb{R}$ let $\mathcal{N}_{z}=\ker(T^{*}-z)$. Then the map
$\Gamma_{0}:\mathcal{N}_{z}\longrightarrow H^{-1/2}(S_{a})$
@@
Surjectivity follows from Lemma 3, by combining the interior and exterior solutions with a common trace $\phi\in H^{1/2}(S_{a})$. More explicitly, if $\phi$ denotes this trace, then the jump of the normal derivative is
$\Gamma_{0}u=[\partial_{r}u]=(\partial_{r}u)^{+}-(\partial_{r}u)^{-}=\big{(}\mathcal{N}_{+}(z)+\mathcal{N}_{-}(z)\big{)}\phi,$
where $\mathcal{N}_{\pm}(z):H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ are the Dirichlet–to–Neumann maps for the interior and exterior domains. It is well known that $\mathcal{N}_{+}(z)+\mathcal{N}_{-}(z)$ is an elliptic pseudodifferential operator of order one, with principal symbol $|\xi|$, and therefore a topological isomorphism $H^{1/2}(S_{a})\to H^{-1/2}(S_{a})$ for nonreal $z$ (cf. Behrndt–Hassi–de Snoo [3, Chapter 6]). This establishes surjectivity. If $u\in\mathcal{N}_{z}$ satisfies $\Gamma_{0}u=0$, then $u^{\pm}$ solve homogeneous Dirichlet problems with nonreal energy, which implies $u^{\pm}\equiv 0$. Hence $\Gamma_{0}$ is also injective. The boundedness of the inverse map follows from the closed graph theorem. $\Box$
Lemma 4 shows that $\Gamma_{0}:\mathcal{N}_{z}\to H^{-1/2}(S_{a})$ is bijective for every nonreal $z$. Since $H^{1/2}(S_{a})$ is infinite dimensional, it follows that the deficiency indices of $T$ are infinite, so $T$ is not essentially self–adjoint. Thus self–adjoint realizations arise precisely from boundary conditions relating $\Gamma_{0}$ and $\Gamma_{1}$. In particular, when the wall strength varies along the surface, $\alpha(\omega)\in L^{\infty}(S_{a};\mathbb{R})$, we introduce the boundary multiplication operator
@@ 4 Spectrum of the penetrable–wall Stark Hamiltonian
Step 3: Mourre estimate and absence of eigenvalues. Consider the pair $(H_{F,0},A_{F})$ in the sense of Mourre theory, where $H_{F,0}$ is the free Stark Hamiltonian and $A_{F}:=-\operatorname{sgn}(F)\,p_{1}$ with $p_{1}=-i\partial_{x_{1}}$ serves as the conjugate operator. A straightforward computation shows that $H_{F,0},iA_{F}]=|F|I$, so the pair $(H_{F,0},A_{F})$ satisfies a strict Mourre estimate on every bounded interval; cf. [9, 11].
Moreover, by Lemma 8 we have that $H_{F,\alpha}-H_{F,0}$ is $H_{F,0}$–compact, and by the Helffer–Sjöstrand functional calculus the spectral projections satisfy
$E_{J}(H_{F,\alpha})-E_{J}(H_{F,0})\ \text{is compact for every bounded interval }J\subset\mathbb{R}.$
Hence, by the general transfer principle for Mourre estimates ([1, Thm. 7.2.9], see also [11, Thm. 2.2 & Thm. 4.3]), the pair $(H_{F,\alpha},A_{F})$ is of class $C^{1}$ on any bounded interval and the strict Mourre estimate
$E_{J}(H_{F,\alpha})[H_{F,\alpha},iA_{F}]E_{J}(H_{F,\alpha})=|F|\,E_{J}(H_{F,\alpha})+K_{J},\qquad K_{J}\text{ compact},$
holds on each bounded $J$. Choosing intervals that cover the real line, one concludes that $H_{F,\alpha}$ has no eigenvalues, and $\sigma(H_{F,\alpha})=\sigma_{\mathrm{ac}}(H_{F,\alpha})=\mathbb{R}$.
$\Box$
@@ 5.1 Complex distortion and analytic continuation
This defines a type–A analytic family of operators in the sense of Kato [12]. The distorted resolvent
@@ 5.2 Distorted boundary operators and Krein formula
and is denoted by $H_{F,\alpha}(\theta)$. For $0<\text{{\rm Im}}\theta<\theta_{0}$, $U_{\theta}$ is boundedly invertible on $L^{2}$ and the distorted free operator $H_{F,0}(\theta)=U_{\theta}H_{F,0}U_{\theta}^{-1}$ is mm–sectorial [8, 9]. We keep the real boundary maps $(\Gamma_{0},\Gamma_{1})$ from Section 2, so the $\delta$–shell boundary condition $\Gamma_{0}=A_{\alpha}^{b}\Gamma_{1}$ remains unchanged under the distortion. What changes is only the Poisson and Weyl operators $\gamma_{F}(\theta;z)$ and $M_{F}(\theta;z)$, which are now constructed using the distorted free resolvent $R_{F,0}(\theta;z)$. Within the quasi boundary triple framework this leads to a closed mm–sectorial extension, whose resolvent is given by the distorted Krein formula (see (5.2)); cf. [3, Sec. 6, Thm. 6.16] and [2, Thm. 2.6]. For $\theta\in\mathbb{R}$ we recover the undistorted self–adjoint extension $H_{F,\alpha}$.
@@
It is important to note that the complex distortion $U_{\theta}$ acts only on the variable in the $x_{1}$–direction of the functions, while the geometric surface $S_{a}=\{x:|x|=a\}$ itself remains fixed. Hence the domains of the boundary maps $(\Gamma_{0},\Gamma_{1})$ and of the multiplication operator $A_{\alpha}^{b}$ are unaffected, and the trace map $H^{1}(\mathbb{R}^{d})\to H^{1/2}(S_{a})$ is stable under distortion. Therefore the quasi boundary triple structure carries over verbatim to the distorted setting.
@@
Here the minus sign is consistent with our convention for the Green’s identity; in the standard formulas for quasi boundary triples the sign depends on the choice of convention, and in our setting the present convention produces the additional factor 1-1.
Step 3 (Ansatz and boundary equation). Let $f\in L^{2}(\mathbb{R}^{d})$ and $z\in\rho(A_{0}(\theta))\cap\rho(A_{\alpha}(\theta))$. Set $u_{0}=(A_{0}(\theta)-z)^{-1}f$, so $(T(\theta)^{*}-z)u_{0}=f$ and $\Gamma_{0}u_{0}=0$. Seek $u=u_{0}-\gamma_{F}(\theta;z)\varphi$ with unknown $\varphi\in H^{-1/2}(S_{a})$. Then
$\Gamma_{0}u=-\varphi,\qquad\Gamma_{1}u=\Gamma_{1}u_{0}-M_{F}(\theta;z)\varphi.$
Imposing the boundary condition $\Gamma_{0}=A_{\alpha}^{b}\Gamma_{1}$ yields
$\varphi\;=\;A_{\alpha}^{b}\bigl{(}\Gamma_{1}u_{0}-M_{F}(\theta;z)\varphi\bigr{)},$
or equivalently
$\bigl{(}I-A_{\alpha}^{b}M_{F}(\theta;z)\bigr{)}\varphi=-\,A_{\alpha}^{b}\Gamma_{1}u_{0},$
from which we obtain
$\varphi=-\,\bigl{(}I-A_{\alpha}^{b}M_{F}(\theta;z)\bigr{)}^{-1}A_{\alpha}^{b}\Gamma_{1}u_{0}.$
Substituting into the ansatz gives
$u\;=\;u_{0}+\gamma_{F}(\theta;z)\,\bigl{(}I-A_{\alpha}^{b}M_{F}(\theta;z)\bigr{)}^{-1}A_{\alpha}^{b}\Gamma_{1}u_{0}.$
$\Box$
@@ 5.3 Boundary characterization of resonances
By the Herbst distortion theory (Theorem 11), if $0<\text{{\rm Im}}~\theta<\theta_{0}$, the discrete eigenvalues of $H_{F,\alpha}(\theta)$ in the lower half-plane are independent of $\theta$. Therefore, the set of resonances of $H_{F,\alpha}$ is well defined and coincides with the zero set of $D_{F}^{(p)}(\theta;z)$ for any admissible $\theta$.
5.4 Quadratic shift and exponential width for the $\ell=0$ bound state
We consider the case of constant wall strength in three dimensions and focus on the spherically symmetric mode. The aim of this subsection is to derive the small–field expansion of the resonance, to obtain the quadratic shift, and to explain the exponential behavior of the resonance width.
At zero field the resolvent kernel is
$G_{0}(x,y;z)=\frac{e^{-\kappa|x-y|}}{4\pi|x-y|},\qquad z=-\kappa^{2}<0,$
and the addition theorem gives
$G_{0}(a\omega,a\omega^{\prime};z)=\frac{\kappa}{4\pi}\sum_{\ell=0}^{\infty}(2\ell+1)i_{\ell}(\kappa a)k_{\ell}(\kappa a)P_{\ell}(\omega\cdot\omega^{\prime}).$
Here $i_{\ell}$ and $k_{\ell}$ are the modified spherical Bessel functions. Hence $M_{0}(z)Y_{\ell m}=\mu_{\ell}(z)Y_{\ell m}$ with
$\mu_{\ell}(z)=\frac{a}{2\kappa}i_{\ell}(\kappa a)k_{\ell}(\kappa a).$
Bound states are determined by $1+\alpha\mu_{\ell}(z)=0$. For $\ell=0$ this relation becomes
$\alpha=-\kappa\left(1+\coth(\kappa a)\right),\qquad E_{0}=-\kappa^{2}<0.$
Small-field expansion for the s-wave ($\ell=0$).
In $d=3$ with constant wall strength $\alpha$, we expand the boundary Weyl operator at small field
$M_{F}(z)=M_{0}(z)-F\,\tau R_{0}(z)x_{1}R_{0}(z)\tau^{*}+F^{2}\,\tau R_{0}(z)x_{1}R_{0}(z)x_{1}R_{0}(z)\tau^{*}+O(F^{3}),$
and seek a resonance branch
$z(F)=E_{0}+Fa_{1}+F^{2}a_{2}+O(F^{3})$
bifurcating from the bound state $E_{0}<0$ at $F=0$. By the selection rule for spherical harmonics, the operator $x_{1}$ couples only spherical harmonics with angular momenta differing by one, $\ell\to\ell\pm 1$. Equivalently, the Gaunt coefficients (or 3j3j–symbols) for the integral $\int_{S^{2}}Y_{\ell m}(\omega)\,\omega_{1}\,Y_{\ell^{\prime}m^{\prime}}(\omega)\,d\omega$ vanish unless $|\ell-\ell^{\prime}|=1$. Therefore the ss–wave ($\ell=0$) component is coupled only to $\ell=1$, and does not couple back to itself. This forces the linear coefficient to vanish, $a_{1}=0$ (see also Appendix A.1 for explicit formulas).
By the selection rule for spherical harmonics ($\ell\to\ell\pm 1$), we have in particular $\langle Y_{00},M_{1}(E_{0})Y_{00}\rangle=0$, so that the linear term vanishes.
Thus the quadratic coefficient takes the explicit form
$a_{2}=-\,\frac{\langle Y_{00},\,M_{2}(E_{0})\,Y_{00}\rangle}{\mu_{0}^{\prime}(E_{0})}\;-\;\frac{\alpha}{\mu_{0}^{\prime}(E_{0})}\sum_{m=-1}^{1}\frac{\big{|}\langle Y_{1m},\,M_{1}(E_{0})\,Y_{00}\rangle\big{|}^{2}}{1+\alpha\,\mu_{1}(E_{0})}.$ (5.3)
Here we implicitly assume that $E_{0}$ is simple and $1+\alpha\,\mu_{1}(E_{0})\neq 0$, i.e. no accidental resonance occurs in the $\ell=1$ channel; otherwise a separate degenerate analysis would be required.
Here $\kappa=\sqrt{-z}$ with $\mathrm{Re}\,\kappa>0$. From
$\mu_{0}(z)=\frac{1-e^{-2\kappa a}}{2\kappa},\qquad(z=-\kappa^{2}),$
we differentiate at $z=E_{0}=-\kappa_{0}^{2}$ to obtain
$\mu_{0}^{\prime}(E_{0})=1-\frac{(1+2\kappa_{0}a)e^{-2\kappa_{0}a}}{4\kappa_{0}^{3}}\;>\;0,$ (5.4)
using $d\kappa/dz=-1/(2\kappa)$.
Turning to the width, the Agmon action along the field direction reads
$S(F)=\int_{a}^{-E_{0}/F}\!\sqrt{Fs-E_{0}}\,ds=\frac{2}{3F}\,(Fa-E_{0})^{3/2}\sim\frac{2}{3F}\,|E_{0}|^{3/2}\qquad(F\downarrow 0),$ (5.5)
hence
$\Gamma(F)\;\asymp\;\exp\!\Big{(}-\,\tfrac{4}{3}\,|E_{0}|^{3/2}/F\Big{)}$
up to a slowly varying prefactor $C(F)\to C_{0}>0$. This type of exponential law is reminiscent of the classical exponential decay law for nuclear $\alpha$–decay [14].
This coincidence of the upper and lower exponential bounds is a classical fact for Stark resonances; see Herbst [9], Harrell–Simon [7], and Wang [16] for rigorous proofs of matching asymptotics.
The prefactor can be derived from the one–dimensional WKB connection at $r=a$ for the $Y_{00}$ component, and is described in the Appendix.
In summary, for the constant wall case and the ss–wave bound state the resonance takes the form
$z(F)=E_{0}+F^{2}a_{2}+O(F^{3})\;-\;i\,C(F)\exp\!\left(-\tfrac{4}{3}\tfrac{|E_{0}|^{3/2}}{F}\right),\qquad F\downarrow 0,$ (5.6)
with $C(F)\to C_{0}>0$ as $F\to 0$. See also Wang [17] for related results on NN–body Stark resonances.